Let's say that a company's profitability, p(t), follows p(t) = t^2 - 5t - 50 with p(t) in units of $1k and t = fiscal months. How long will it take the company to be $100k in the black (i.e., p(t) = $100,000)?
a. 14 years
b. 1.5 years
c. 1.25 years
d. 6 months

Answer by edjones(8007)

p(t) = t^2 - 5t - 50
t^2 - 5t - 50 =100
t^2-5t-150=0
(t-15)(t+10)=0
t=15 fiscal months
.
Ed